# Course Allocation with Credits via Stable Matching

### v1

@@ 5.2 With lower quotas
As $M^{\prime}$ is student-complete, every student in $P$ is matched to some $b_{i}^{l}$, where $l=1$ or $l=2$. In order to satisfy the lower quota of $b_{i}^{l}$ and make sure $M^{\prime}$ is pair-stable, $(s_{i}^{l},b_{i}^{l})\in M^{\prime}$ too. As $M^{\prime}$ is student-complete, $s_{i}^{l}$ is matched to some $c_{j}$ and $M^{\prime}$ includes the pair $(s_{i}^{4},d_{i}^{1})$. As the lower quota of $d_{i}^{1}$ is $2$, $M^{\prime}$ must include the pair $(s_{i}^{m},d_{i}^{l})$, where $m=1$ if $l=2$ and vice versa. This implies that $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$. Hence, for all $1\leq i\leq n_{2}$, if $s_{i}^{1}$ is matched to some $c_{j}$ then $s_{i}^{2}$ is matched to $d_{i}^{1}$, and if $s_{i}^{2}$ is matched to some $c_{k}$ then $s_{i}^{1}$ is matched to $d_{i}^{1}$. Thus, we can build a matching $M$ in $G$ of size $K$ by taking every pair of the form $(s_{i}^{l},c_{j})\in M$, of which there are $K$, and adding the pair $(u_{j},w_{i})$ to $M$. We now show that $M$ is a maximal matching in $G$.
@@
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $c_{j}$ is unmatched and $s_{i}^{3}$ is matched with $e$. However, $s_{i}^{3}$ prefers $c_{j}$ over $e$ and $c_{j}$ prefers $s_{i}^{3}$ over being unmatched. Thus, $(s_{i}^{3},c_{j})$ is a blocking pair, a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
Hence, there exists a maximal matching of size $K$ in $G$ if and only if there exists a student-complete first-coalition-stable matching in $I$, respecting lower quotas of every open course. As the first problem is NP-complete, so is the second one. ∎
@@ 6 Conclusion
We have introduced a new model of the Course Allocation problem, which allows for courses with different numbers of credits and constraints such as courses that run at overlapping times. We have considered four natural definitions of stability in this context, and for each one we studied the problem of determining whether a matching is stable and finding a stable matching if one exists. Additionally, we studied the problem with master lists and lower quotas, and considered the problems of finding a stable matching and a maximum size stable matching.
The main problem remaining is finding a coalition-stable matching under arbitrary preferences, and determining whether one always exists. Another possible research direction is studying stability under additively separable preferences, where students assign a utility to a course and prefer subsets of courses where the sum of utilities is higher cechlarova_pareto_2018 . A restriction of additively separable preferences is lexicographic preferences, where students compare two bundles of courses by first checking the first course of each bundle, then the second course, and so on cechlarova_pareto_2018 . If we also require that students participating in a blocking (according to additively separable or lexicographic preferences) coalition do not reduce their total number of credits, then these definitions are more strict than first-coalition stability but less so than coalition stability.
Other possibilities include finding stable matchings with the lowest number of blocking pairs/coalitions when there is no stable matching, approximation algorithms for finding a maximum size stable matching, or modifying course upper quotas to ensure the existence of a stable matching, following the work of Nguyen and Vohra nguyen_2018_near .

### v2

@@ 5.2 With lower quotas
As $M^{\prime}$ is student-complete, every student $p_{k}\in P$ (there are $K$ such students) is matched to some $b_{i}^{r}\in B$, where $r\in\{1,2\}$. In order to satisfy the lower quota of $b_{i}^{r}$, and by the pair stability of $M^{\prime}$, $(s_{i}^{r},b_{i}^{r})\in M^{\prime}$ also. As $M^{\prime}$ is student-complete, $(s_{i}^{r},c_{j})\in M^{\prime}$ for some $c_{j}\in C^{\prime}$. As the lower quota of $d_{i}^{1}$ is $2$, and we have already established that $d_{i}^{1}$ is open, $M^{\prime}$ must include the pair $(s_{i}^{3-r},d_{i}^{1})$. This implies that $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ by pair stability. Hence, for each $i$ ($1\leq i\leq n_{2}$), if $(s_{i}^{1},c_{j})\in M$ for some $c_{j}\in C^{\prime}$ then $(s_{i}^{2},d_{i}^{1})\in M^{\prime}$, whilst if $(s_{i}^{2},c_{k})\in M^{\prime}$ for some $c_{k}\in C^{\prime}$ then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$. Thus, we can build a matching $M$ in $G$ by adding the pair $(u_{j},w_{i})$ to $M$ for every pair of the form $(s_{i}^{r},c_{j})\in M^{\prime}$. It follows by the fact that $|S_{M}(e)|=n_{2}-K$ that $|M|=K$. We now show that $M$ is a maximal matching in $G$.
@@
Looking for a contradiction, assume that $M$ is not maximal, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other in $G$. This implies that in $M^{\prime}$, $c_{j}$ is unmatched and $s_{i}^{3}$ is matched with $e$. However, $s_{i}^{3}$ prefers $c_{j}$ over $e$, whilst $c_{j}$ is unmatched, has lower quota 0, and finds $s_{i}^{3}$ acceptable. Thus, $(s_{i}^{3},c_{j})$ is a blocking pair of $M^{\prime}$, a contradiction to the pair stability of $M^{\prime}$ in $I$. Therefore, $M$ is maximal in $G$. ∎
Hence, there exists a maximal matching of size $K$ in $G$ if and only if there exists a student-complete first-coalition-stable matching in $I$, respecting lower quotas of every open course. As exact-mm is NP-complete, it follows that ca-p-ml-lq-cl-max is NP-hard. ∎
@@ 6 Conclusion
We have introduced a new model of the course allocation problem, which allows for courses with different numbers of credits and constraints such as courses that run at overlapping times. We have considered four natural definitions of stability in this context, and for each one we studied the problem of determining whether a matching is stable and finding a stable matching if one exists. Additionally, we studied the problem with master lists, and considered the problems of finding a stable matching and a maximum size stable matching, both in the absence of, and in the presence of, lower quotas.
The main open problem remaining is to establish the computational complexity of finding a coalition-stable matching under arbitrary preferences, and indeed to determine whether one always exists. Another possible research direction is to study stability under additively separable preferences, where students assign a utility to a course and prefer subsets of courses where the sum of utilities is higher Cechlárová et al. (2018). A restriction of additively separable preferences is lexicographic preferences, where students compare two bundles of courses by first checking the first course of each bundle, then the second course, and so on Cechlárová et al. (2018). If we also require that students participating in a blocking (according to additively separable or lexicographic preferences) coalition do not reduce their total number of credits, then these definitions of stability are less strict than first-coalition stability but more so than coalition stability.
Other possible future research directions include considering problems involving finding stable matchings with the minimum number of blocking pairs/coalitions when there is no stable matching, developing approximation algorithms for finding a maximum size stable matching, such as the 3-approximation algorithm for the problem of finding a maximum size occupancy-stable (equivalent to pair-size-stable) matching by Balasundaram et al. Balasundaram et al. (2025), or investigating problems involving modifying course upper quotas to ensure the existence of a stable matching, following the work of Nguyen and Vohra Nguyen and Vohra (2018).
Acknowledgements
The authors would like to thank Zoe Graves and Matt Wierzbicki of the School of Law, University of Glasgow, for inspiring the ideas behind our model for ca; Mathijs Barkel and Fraser Paterson for valuable discussions concerning algorithms for ca-c-ml-max; and the SAGT reviewers for their valuable comments and helpful suggestions. José Rodríguez is supported by a studentship from the University of Glasgow, College of Science and Engineering. The authors have no competing interests to declare that are relevant to the content of this article. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.